# Distilling Morphology-Conditioned Hypernetworks for Efficient Universal Morphology Control

### v1

@@ Abstract
Learning a universal policy across different robot morphologies can significantly improve learning efficiency and enable zero-shot generalization to unseen morphologies. However, learning a highly performant universal policy requires sophisticated architectures like transformers (TF) that have larger memory and computational cost than simpler multi-layer perceptrons (MLP). To achieve both good performance like TF and high efficiency like MLP at inference time, we propose HyperDistill, which consists of: (1) A morphology-conditioned hypernetwork (HN) that generates robot-wise MLP policies, and (2) A policy distillation approach that is essential for successful training. We show that on UNIMAL, a benchmark with hundreds of diverse morphologies, HyperDistill performs as well as a universal TF teacher policy on both training and unseen test robots, but reduces model size by 6-14 times, and computational cost by 67-160 times in different environments. Our analysis attributes the efficiency advantage of HyperDistill at inference time to knowledge decoupling, i.e., the ability to decouple inter-task and intra-task knowledge, a general principle that could also be applied to improve inference efficiency in other domains.
@@ 2.2 Architectures for Universal Morphology Control
When learning a universal MLP policy, to handle the variable number of limbs across different robots, it is common to assume a maximal limb number $N_{\text{max}}$, and zero-pad the state and action vectors to this maximal length so that the state and action dimensions are aligned across different robots. As the universal policy also conditions on the morphology context, the context features $c_{k}^{i}$ are usually concatenated with the state features $s_{k,t}^{i}$ as node-wise input (Gupta et al., 2022). So the final input to the MLP policy is $x=\left[x_{1},x_{2},\dots,x_{N_{k}},\vec{0},\dots,\vec{0}\right]$, where $x_{i}=[s_{i},c_{i}]$, and we omit the robot index $k$ and time index $t$ thereafter for simplicity of notation. For the input layer of the MLP, we have $h^{(0)}=\sigma\left(W^{\text{in}}x+b^{\text{in}}\right)=\sigma\left(\sum_{i=1}^{N_{k}}W^{\text{in}}_{i}x_{i}+b^{\text{in}}\right)$, where $\sigma$ is the activation function, $W_{i}^{\text{in}}$ is the subset of the weight matrix that encodes the input features of node $i$. For the $l$-th hidden layer, we have $h^{(l+1)}=\sigma\left(W^{(l)}h^{(l)}+b^{(l)}\right)$. For the output layer, we have $a=W^{\text{out}}h^{(L)}+b^{\text{out}}$, which can be decomposed into node-wise action $a_{i}=W_{i}^{\text{out}}h^{(L)}+b_{i}^{\text{out}}$.
Unlike MLP, GNN (Wu et al., 2020) and TF (Vaswani et al., 2017) can naturally handle variable numbers of homogeneous elements. GNN can be directly applied to different morphology graphs, while TF treats each limb as a token. Unlike in many domains where tokens are processed sequentially, in universal morphology control, TF processes different nodes in parallel to model their spatial interactions.
@@ 3 HyperDistill
This section introduces HyperDistill, which achieves both good performance and high efficiency at inference time. In Section 3.1, we analyze the limitations of existing methods for universal morphology control, which motivates us to introduce HN as a solution in Section 3.2. In Section 3.3, we discuss how and why we train the HN via policy distillation, and analyze some of its critical algorithmic features.
@@ 3.1 Limitations of Existing Architectures
A further issue with MLP is that it is not invariant to the order of limbs across robots. TF is order-invariant, as all the limbs across different robots share the same set of parameters. By contrast, in an MLP, only the limbs with the same index across different robots share the same subset of parameters in the input and output layers, so how we order the limbs influences the policy output. As we do not have a consistent way to order the limbs across different morphologies, MLP can overfit to spurious patterns in the manually chosen limb indexing method, harming generalization.
@@ 3.2 HyperDistill Architecture
The overall architecture is shown in Figure 1. Unlike prior work in HN, a unique challenge in applying HN to universal morphology control is that the input and output dimensions of the MLP base network vary across robots due to the variable number of limbs. We tackle this challenge by generating limb-wise parameters for the input and output layers, as the subset of parameters corresponding to each limb still has fixed dimension based on the modular space assumption in Section 2.1. Specifically, for the input layer, we have $h^{(0)}=\sigma\left(\sum_{i}W^{\text{in}}_{i}x_{i}+b^{\text{in}}_{i}\right)$, where $W_{i}^{\text{in}}=\text{HN}_{W}^{\text{in}}\left(f\left(c_{i}\right)\right)$, and $b^{\text{in}}_{i}=\text{HN}_{b}^{\text{in}}\left(f(c_{i})\right)$. As the whole policy conditions on the morphology context through HN, we no longer need to concatenate context features $c_{i}$ to the network input, so we have $x_{i}=s_{i}$. Similarly, for the output layer, we have $a_{i}=W_{i}^{\text{out}}h^{(L)}+b_{i}^{\text{out}}$, where $W_{i}^{\text{out}}=\text{HN}_{W}^{\text{out}}\left(f(c_{i})\right)$, and $b_{i}^{\text{out}}=\text{HN}_{b}^{\text{out}}\left(f(c_{i})\right)$. For the hidden layers, we first aggregate the context embedding of different limbs using the mean to get a context embedding for the whole morphology as $e_{m}=\frac{1}{N}\sum_{i}e_{i}$, then pass $e_{m}$ through HN output heads to generate the hidden layer parameters, i.e., $h^{(l+1)}=\sigma\left(W^{(l)}h^{(l)}+b^{(l)}\right)$, where $W^{(l)}=\text{HN}_{W}^{(l)}(e_{m})$, and $b^{(l)}=\text{HN}_{b}^{(l)}(e_{m})$.
@@ 3.3 HyperDistill Training via Policy Distillation
Instead, we adopt policy distillation (Parisotto et al., 2015; Rusu et al., 2015) by first training a universal TF policy, then distilling it into an HN policy via behavior cloning, which replaces RL training with more stable supervised learning to alleviate the optimization challenge. Moreover, as BC empirically requires much less time to train than RL, we can reuse the same pre-trained policy for more efficient evaluation of different algorithmic choices.
@@
Regularization is a common approach to reduce overfitting and improve generalization (Cobbe et al., 2019). We consider it to be especially important for HyperDistill, as HN may be more prone to overfitting than other architectures. In prior work, morphology context is usually just used as an additional policy input (Gupta et al., 2022), while in HyperDistill, it is used as HN input to generate the whole control policy. As we only have a few hundred different robots for training, which is much less than the number of parameters in the HN, there is a high chance of overfitting. To reduce overfitting, we apply dropout to the context embedding $e_{i}$ and $e_{m}$. This can be seen as regularization in task space, which encourages the HN to learn an ensemble of different MLP policies that can all work well on the same robot. It can also be seen as domain randomization (Tobin et al., 2017) by making the generated policy more robust to changes in morphology context. We leave it for future work to investigate other regularization methods like weight decay.
@@ 4.1 Experimental Setup
Teacher Policy and Data Collection
For each environment, we train a universal TF policy on the 100 training robots as the teacher policy. We adopt ModuMorph (Xiong et al., 2023) as the TF architecture, as it achieves SOTA performance on the UNIMAL benchmark. To collect data for policy distillation, we first generate an augmented robot set of 1,000 PD robots by mutating the 100 training robots. See Appendix B.1 for how we do the mutation. Then we collect 8,000 transitions from each PD robot, forming a multi-robot dataset with 8M transitions for policy distillation.
@@
We compare HyperDistill with the following architectures as the student policy: (1) TF (oracle): A TF with the same architecture as the teacher. It serves as a performance upper bound on how well the student can perform without architecture misalignment. (2) TF (compressed): A TF with a similar number of parameters as the base MLP in HyperDistill. Nonetheless, the computational cost of TF (compressed) is still higher due to parameter sharing. (3) MLP: A multi-robot MLP with the same architecture as the base MLP in HyperDistill, which is used to validate the advantages of HyperDistill over MLP as discussed in Section 3.2. We also compare with training an HN policy via RL (HN-RL) to show the importance of distillation.
@@
The distillation process runs for 150 epochs, with a mini batch size of 5120. We use Adam with a learning rate of 0.0003, and clip the gradient norm to 0.5. We run three random seeds for each method in each environment, and report the average performance and standard deviation. For HyperDistill, we apply dropout to context embedding with $p=0.1$. See Appendix B.3 for the size of each student model in each environment.  [^1] Code will be publicly available upon acceptance.
@@ 4.2 Main Results
Figures 2 illustrates how different methods generalize to test robots during the distillation process. Table 1 compares the model size and FLOPs of different methods at inference time. Due to space limitations, see Figure 6 in the appendix for different methods’ performance on the training robots.
@@
FT TF (oracle) 1.73 M 14.0 39.86 M 160.8
TF (compressed) 0.15 M 1.2 1.78 M 7.2
MLP 0.23 M 1.9 0.46 M 1.9
@@
VT TF (oracle) 1.97 M 6.6 40.33 M 67.2
TF (compressed) 0.31 M 1.0 3.58 M 6.0
MLP 0.41 M 1.4 0.82 M 1.4
@@
Obstacle TF (oracle) 2.02 M 6.7 40.43 M 67.4
TF (compressed) 0.31 M 1.0 3.58 M 6.0
MLP 0.41 M 1.4 0.82 M 1.4
@@
Table 1: Model size and FLOPs of different methods at inference time. Abs. denotes the absolute value, and Rel. denotes the relative value w.r.t. HyperDistill. See Appendix B.2 for how we compute the FLOPs, and C.2 for more analysis on the results in this table.
HyperDistill achieves performance similar to TF (oracle), matching the TF teacher on both the training and test robots in all the three environments, while reducing model size by 6-14 times, FLOPs by 67-160 times in different environments. Furthermore, TF (compressed) cannot match the performance of HyperDistill in all the three environments, proving that the TF architecture needs to trade off between performance and efficiency, while HyperDistill can achieve both at inference time. HyperDistill also significantly outperforms MLP, which validates the importance of introducing the HN. Moreover, although TF (compressed) and MLP have similar model size, TF (compressed) still outperforms MLP, and their performance gap on the test robots is larger than the gap on the training robots, which may be overfitting caused by the order-invariance issue of MLP as discussed in Section 3.1.
@@ 4.3 Ablation Studies
In this section, we investigate how the algorithmic choices in PD influence generalization performance of the student. Due to space limitations, see Appendix C.3 for ablations on architecture choices in HyperDistill. To save computation, we run PD for only 50 epochs as most methods can already converge within this budget, and experiment only in FT unless otherwise stated.
@@ 6 Related Work
While generalization across robots with variations only in kinematics or dynamics parameters has been extensively studied in prior work (Tobin et al., 2017; Peng et al., 2018; Clavera et al., 2019; Ghadirzadeh et al., 2021; Feng et al., 2022), universal morphology control poses a more challenging task of generalization across morphologies. While learning an MLP policy with zero-padding is feasible to work across morphologies, it is significantly outperformed by GNN (Wang et al., 2018; Pathak et al., 2019; Huang et al., 2020) and TF (Kurin et al., 2020; Dong et al., 2022; Furuta et al., 2022; Gupta et al., 2022; Hong et al., 2022; Trabucco et al., 2022; Chen et al., 2023; Xiong et al., 2023). Furthermore, TF outperforms GNN by better modeling the interactions between distant limbs in a robot (Kurin et al., 2020), thus has been adopted by most recent works. However, these more sophisticated models introduce higher computational cost during deployment, which we aim to tackle by generating a compact MLP controller at inference time with morphology-conditioned HN. Similar to our work, ModuMorph (Xiong et al., 2023) also utilizes HN. However, the main body of ModuMorph is still a TF, while HN is only used to generate node-wise embedding and output layers, thus cannot reduce computational cost. By contrast, we use HN to generate an MLP controller which is much more efficient at inference time.

### v2

@@ Abstract
Learning a universal policy across different robot morphologies can significantly improve learning efficiency and enable zero-shot generalization to unseen morphologies. However, learning a highly performant universal policy requires sophisticated architectures like transformers (TF) that have larger memory and computational cost than simpler multi-layer perceptrons (MLP). To achieve both good performance like TF and high efficiency like MLP at inference time, we propose HyperDistill, which consists of: (1) A morphology-conditioned hypernetwork (HN) that generates robot-wise MLP policies, and (2) A policy distillation approach that is essential for successful training of the HN. We show that on UNIMAL, a benchmark with hundreds of diverse morphologies, HyperDistill performs as well as a universal TF teacher policy on both training and unseen test robots, but reduces model size by 6-14 times, and computational cost by 67-160 times in different environments. Our analysis attributes the efficiency advantage of HyperDistill at inference time to knowledge decoupling, i.e., the ability to decouple inter-task and intra-task knowledge, a general principle that could also be applied to improve inference efficiency in other domains. The code is publicly available at https://github.com/MasterXiong/Universal-Morphology-Control.
@@ 2.2 Architectures for Universal Morphology Control
When learning a multi-robot MLP policy, to handle the variable number of limbs across different robots, it is common to assume a maximal limb number $N_{\text{max}}$, and zero-pad the state and action vectors to this maximal length so that the state and action dimensions are aligned across different robots. As the universal policy also conditions on the morphology context, the context features $c_{k}^{i}$ are usually concatenated with the state features $s_{k,t}^{i}$ as node-wise input (Gupta et al., 2022). So the final input to the multi-robot MLP policy is $x=\left[x_{1},x_{2},\dots,x_{N_{k}},\vec{0},\dots,\vec{0}\right]$, where $x_{i}=[s_{i},c_{i}]$, and we omit the robot index $k$ and time index $t$ thereafter for simplicity of notation. For the input layer of the MLP, we have $h^{(0)}=\sigma\left(W^{\text{in}}x+b^{\text{in}}\right)=\sigma\left(\sum_{i=1}^{N_{k}}W^{\text{in}}_{i}x_{i}+b^{\text{in}}\right)$, where $\sigma$ is the activation function, $W_{i}^{\text{in}}$ is the subset of the weight matrix that encodes the input features of node $i$. For the $l$-th hidden layer, we have $h^{(l+1)}=\sigma\left(W^{(l)}h^{(l)}+b^{(l)}\right)$. For the output layer, we have $a=W^{\text{out}}h^{(L)}+b^{\text{out}}$, which can be decomposed into node-wise action $a_{i}=W_{i}^{\text{out}}h^{(L)}+b_{i}^{\text{out}}$.
Unlike MLP, GNN (Wu et al., 2020) and TF (Vaswani et al., 2017) can naturally handle variable numbers of homogeneous elements. GNN can be directly applied to different morphology graphs, while TF treats each limb as a token. Unlike in many domains where tokens are processed sequentially, in universal morphology control, TF processes different nodes in parallel to model their spatial interactions. A typical TF architecture for universal morphology control consists of a linear embedding layer that embeds node-wise input features, multiple attention blocks that model limb interactions, and an MLP decoder that maps the node-wise embedding outputed by the attention blocks to node-wise actions (Gupta et al., 2022; Xiong et al., 2023).
@@ 3 HyperDistill
This section introduces HyperDistill, which achieves both good performance and high efficiency at inference time. In Section 3.1, we analyze the limitations of existing methods for universal morphology control, which motivates us to introduce HN as a solution in Section 3.2. In Section 3.3, we discuss why and how we train the HN via policy distillation, and analyze some of its critical algorithmic designs.
@@ 3.1 Limitations of Existing Architectures
A further issue with MLP is that it is not invariant to the order of limbs across robots. TF is order-invariant, as all the limbs across different robots share the same set of parameters. By contrast, in a multi-robot MLP, only the limbs with the same index across different robots share the same subset of parameters in the input and output layers, so how we order the limbs influences the policy output. As we do not have a consistent way to order the limbs across different morphologies, multi-robot MLP can overfit to spurious patterns in the manually chosen limb indexing method, harming generalization.
@@ 3.2 HyperDistill Architecture
The overall architecture is shown in Figure 1. Unlike prior work in HN, a unique challenge in applying HN to universal morphology control is that the input and output dimensions of the MLP base network vary across robots due to the variable number of limbs. We tackle this challenge by generating limb-wise parameters for the input and output layers, as the subset of parameters corresponding to each limb still has fixed dimension based on the modular space assumption in Section 2.1. Specifically, for the input layer, we have $h^{(0)}=\sigma\left(\sum_{i}W^{\text{in}}_{i}x_{i}+b^{\text{in}}_{i}\right)$, where $W_{i}^{\text{in}}=\text{HN}_{W}^{\text{in}}(e_{i})$, $b^{\text{in}}_{i}=\text{HN}_{b}^{\text{in}}(e_{i})$, and $e_{i}=f(c_{i})$ is the context embedding of limb $i$ generated by the context encoder $f$. As the whole policy conditions on the morphology context through HN, we no longer need to concatenate context features $c_{i}$ to the network input, so we have $x_{i}=s_{i}$. Similarly, for the output layer, we have $a_{i}=W_{i}^{\text{out}}h^{(L)}+b_{i}^{\text{out}}$, where $W_{i}^{\text{out}}=\text{HN}_{W}^{\text{out}}(e_{i})$, and $b_{i}^{\text{out}}=\text{HN}_{b}^{\text{out}}(e_{i})$. For the hidden layers, we first aggregate the context embedding of different limbs using the mean to get a context embedding for the whole morphology as $e_{m}=\frac{1}{N}\sum_{i}e_{i}$, then pass $e_{m}$ through HN output heads to generate the hidden layer parameters, i.e., $h^{(l+1)}=\sigma\left(W^{(l)}h^{(l)}+b^{(l)}\right)$, where $W^{(l)}=\text{HN}_{W}^{(l)}(e_{m})$, and $b^{(l)}=\text{HN}_{b}^{(l)}(e_{m})$.
@@ 3.3 HyperDistill Training via Policy Distillation
Instead, we adopt policy distillation (Parisotto et al., 2015; Rusu et al., 2015) by first training a universal TF policy, then distilling it into an HN policy via behavior cloning, which replaces RL training with more stable supervised learning to alleviate the optimization challenge. Moreover, as BC empirically requires much less time to train than RL, we can reuse the same pre-trained teacher policy for more efficient evaluation of different algorithmic choices.
@@
Regularization is a common approach to reduce overfitting and improve generalization (Cobbe et al., 2019). We consider it to be especially important for HyperDistill, as HN may be more prone to overfitting than other architectures. In prior work, morphology context is usually just used as an additional policy input (Gupta et al., 2022), while in HyperDistill, it is used as HN input to generate the whole control policy. As we only have a few hundred different robots for PD, which is much less than the number of parameters in the HN, there is a high chance of overfitting. To reduce overfitting, we apply dropout to the context embedding $e_{i}$ and $e_{m}$. This can be seen as regularization in task space, which encourages the HN to learn an ensemble of different MLP policies that can all work well on the same robot. It can also be seen as domain randomization (Tobin et al., 2017) by making the generated policy more robust to changes in morphology context. We leave it for future work to investigate other regularization methods like weight decay.
@@ 4.1 Experimental Setup
Teacher Policy
For each environment, we train a universal TF policy on the 100 training robots as the teacher policy. We adopt ModuMorph (Xiong et al., 2023) as the TF teacher, as it achieves SOTA performance on the UNIMAL benchmark. ModuMorph differs from a standard TF architecture in two ways: (1) In the attention blocks, it computes a fixed attention matrix solely conditioned on the morphology context, which performs better than computing dynamic attention weights based on limb observations. (2) The embedding layer and the MLP decoder are generated by HNs in ModuMorph. Although ModuMorph and HyperDistill both use HNs, their motivations are significantly different. ModuMorph adopts HNs to better model the diverse behaviors across limbs, but is still a large TF-based model with high inference costs. By contrast, HyperDistill utilizes HNs to enable knowledge decoupling for efficient inference. This novel perspective on the role of HNs is a key contribution of our method. Furthermore, as we show in Section 4.2, a TF-based model can be compressed to a much smaller size without much performance loss, but only when equipped with these HN-generated layers, while a standard TF cannot.
Data Collection for Distillation
To collect data for policy distillation, we first generate an augmented robot set of 1,000 PD robots by mutating the 100 training robots. See Appendix B.1 for how we do the mutation. Then we collect 8,000 transitions from each PD robot, forming a multi-robot dataset with 8M transitions for policy distillation.
@@
We compare HyperDistill with the following architectures as the student policy: (1) ModuMorph (oracle): A ModuMorph student with the same architecture as the teacher. It serves as a performance upper bound on how well the student can perform without architecture misalignment. (2) TF (compressed): A standard TF with a similar number of parameters as the base MLP in HyperDistill. It is used to validate that standard TF cannot achieve both good performance and high efficiency at the same time like HyperDistill. (3) ModuMorph (compressed): A ModuMorph student with a similar number of parameters as the base MLP in HyperDistill. As it adopts HNs to generate some layers, we expect it to have better knowledge decoupling ability than standard TF, but may still be less efficient than HyperDistill due to the attention blocks. (4) Multi-robot MLP: A multi-robot MLP with the same architecture as the base MLP in HyperDistill, which is used to validate the advantages of HyperDistill over MLP as discussed in Section 3.2. Finally, we also compare with training an HN policy via RL (HN-RL) to show the importance of distillation.
@@
The distillation process runs for 150 epochs, with a mini batch size of 5120. We use Adam with a learning rate of 0.0003, and clip the gradient norm to 0.5. We run three random seeds for each method in each environment, and report the average performance with standard error. For HyperDistill, we apply dropout to context embedding with $p=0.1$. See Appendix B.3 for the size of each student model in each environment.
@@ 4.2 Main Results
Figures 2 and 3 illustrate how different methods perform on the training and test robots during the distillation process. Table 1 compares the model size and FLOPs of different methods at inference time.
@@
FT ModuMorph (oracle) 1.73 M 14.0 39.86 M 160.8
TF (compressed) 0.14 M 1.1 3.39 M 13.7
ModuMorph (compressed) 0.15 M 1.2 1.78 M 7.2
Multi-robot MLP 0.23 M 1.9 0.46 M 1.9
@@
VT ModuMorph (oracle) 1.97 M 6.6 40.33 M 67.2
TF (compressed) 0.31 M 1.0 5.51 M 9.2
ModuMorph (compressed) 0.39 M 1.3 2.26 M 3.8
Multi-robot MLP 0.41 M 1.4 0.82 M 1.4
@@
Obstacle ModuMorph (oracle) 2.02 M 6.7 40.43 M 67.4
TF (compressed) 0.32 M 1.0 5.61 M 9.3
ModuMorph (compressed) 0.44 M 1.5 2.35 M 3.9
Multi-robot MLP 0.41 M 1.4 0.82 M 1.4
@@
Table 1: Model size and FLOPs of different methods at inference time. Abs. denotes the absolute value, and Rel. denotes the relative value w.r.t. HyperDistill. See Appendix B.2 for how we compute the FLOPs, and C.1 for more analysis on the results in this table.
HyperDistill achieves performance similar to ModuMorph (oracle), matching the teacher’s performance on both the training and test robots in all the three environments, while reducing model size by 6-14 times, FLOPs by 67-160 times in different environments.
TF (compressed) cannot match the performance of HyperDistill in all the three environments, as standard TF needs to trade off between performance and efficiency due to a lack of knowledge decoupling.
As expected, ModuMorph (compressed) consistently outperforms TF (compressed), as it generates some layers of the TF via HNs, which enables better knowledge decoupling. However, it still lags behind HyperDistill w.r.t. both generalization performance and efficiency, as the knowledge encoded in the attention blocks still cannot be decoupled. This result further indicates that, in contrast to previous work (Kurin et al., 2020; Gupta et al., 2022; Xiong et al., 2023), we may not need complicated attention modules to achieve good performance for universal morphology control. In addition, the performance gap between ModuMorph (compressed) and HyperDistill is larger in FT than in VT and Obstacle, possibly because for the latter two environments, ModuMorph has more layers in the HN-generated decoder, which makes it more similar to HyperDistill.
HyperDistill also significantly outperforms multi-robot MLP, which validates the importance of the HN. Moreover, the performance gap between multi-robot MLP and other methods is much larger on the test robots than on the training ones, which may be overfitting due to the order-invariance issue of MLP discussed in Section 3.1.
@@ 4.3 Ablation Studies
In this section, we investigate how the algorithmic choices in PD influence generalization performance of the student. Due to space limitations, see Appendix C.2 for ablations on architecture choices in HyperDistill. To save computation, we run PD for only 50 epochs as most methods can already converge within this budget, and experiment only in the FT experiment unless otherwise stated.
@@ 6 Related Work
While generalization across robots with variations only in kinematics or dynamics parameters has been extensively studied in prior work (Tobin et al., 2017; Peng et al., 2018; Clavera et al., 2019; Ghadirzadeh et al., 2021; Feng et al., 2022), universal morphology control poses a more challenging task of generalization across morphologies. While learning an MLP policy with zero-padding is feasible to work across morphologies, it is significantly outperformed by GNN (Wang et al., 2018; Pathak et al., 2019; Huang et al., 2020) and TF (Kurin et al., 2020; Dong et al., 2022; Furuta et al., 2022; Gupta et al., 2022; Hong et al., 2022; Trabucco et al., 2022; Chen et al., 2023; Xiong et al., 2023). Furthermore, TF outperforms GNN by better modeling the interactions between distant limbs in a robot (Kurin et al., 2020), thus has been adopted by most recent works. However, these more sophisticated models introduce higher computational cost during deployment, which we aim to tackle by generating a compact MLP controller at inference time with morphology-conditioned HN. Similar to our work, ModuMorph (Xiong et al., 2023) also utilizes HN. However, ModuMorph introduces HNs to better model the diverse behaviors across different limbs and only uses HNs to generate embedding and decoder layers inside a TF, while our method introduces HNs to improve inference efficiency via knowledge decoupling and generate a whole MLP controller with HNs.